OpenStudy (anonymous):

Write the equation of a line parallel to the line 2x + 3y = 5 and passes through the point (9, -3)

OpenStudy (solomonzelman):

I will give you a similar example, OK? http://openstudy.com/study#/updates/52697f82e4b040e33d2f88e7

OpenStudy (anonymous):

Okay give me a moment to look over the example. : )

OpenStudy (solomonzelman):

that one is about perpendicular line not parallel, do you want to do this one together?

OpenStudy (solomonzelman):

1) put ' 2x + 3y = 5 ' inot a standard from (y=mx+b)

OpenStudy (solomonzelman):

into, not inot

OpenStudy (anonymous):

Umm.. ' 3y= -2x+5 '

OpenStudy (solomonzelman):

OK, now divide all sides by 3, to set it equal to y

OpenStudy (anonymous):

' y = -2/3x + 5/3 '

OpenStudy (solomonzelman):

yep!

OpenStudy (solomonzelman):

now what point does your parallel line pass through? plug in this point into the above equation ' y = -2/3x + 5/3 ' the slope is the same, don't touch it, b/c the lines are parallel.

OpenStudy (solomonzelman):

I mean y = -2/3x + b plug in your coordinate to find b

OpenStudy (anonymous):

So I plug -3 into b? because -3 is the point in which it crosses the y-axis? Or am I all confused o~o I kinda suck at understanding this stuff.. Sorry

OpenStudy (solomonzelman):

no plug in ' (9, -3)' into y = -2/3x + b to find the b.

OpenStudy (anonymous):

I don't know where to plug it in at ;-;

OpenStudy (solomonzelman):

what is the x coordinate in (9, -3)? and what is the y coordinate in (9, -3)?

OpenStudy (anonymous):

9=x -3=y

OpenStudy (solomonzelman):

plug that into y = -2/3x + b

OpenStudy (solomonzelman):

plug 9 for x, and -3 for y into y = -2/3x + b

OpenStudy (anonymous):

I was showing my friend.. Just about to say I get it now.. lol I'm slow sometimes.. -3 = -2/3(9) + b

OpenStudy (solomonzelman):

so what is the b?

OpenStudy (solomonzelman):